OpenStudy (anonymous):

3. How many solutions does the system of equations have? x=-4y + 4 and 2x + 8y=8 A.one B.two C.infinitely many D. none 4. how many solutions? y=5x+7 and 3y-15x=18 one two infinitely many none

OpenStudy (kc_kennylau):

Express all of 'em into slope-intercept form, if only the slopes are the same, no solution. If the slopes AND the y-intercepts are the same, infinitely many solution. Otherwise, one solution.

OpenStudy (anonymous):

5. how many solutions? x - 2y=6 and 3x-6y=18 one two many none 6.how may solutions? y-6x=-3 and 4y-24x=-16

OpenStudy (kc_kennylau):



OpenStudy (anonymous):

infinity

OpenStudy (anonymous):

for which one